When 358, 528 and 698 are in turn divided by a certain number, the remainder is 1 in each case. Find the number.-
Answer by mananth(16949)

When 358, 528 and 698 are in turn divided by a certain number, the remainder is 1 in each case. Find the number.-
..
the difference between the numbers is 170
528-358=170
698-528=170
..
170= 17*10
!7 is the number